Q1 Develop the series solutions of the following differential equations by Frobenius Method (1) **+2 y + x y = 0 xy + y =
Answer #1

| Page No. so ntry 8 n=o nto-2 x = intr-1) Intro y antr- gines the indreal conation Given xyl +2y + xy=0 is regular singularPage Noi 8=0,-) nave from (3) we recursive formula. following Intr + 1) (n+8+2) Entl teng 1 Enti Cn-l (n+&+1) (n+8+2) we havePage No: Now for x2 x4 + 120 21 2 x4 = 4, (n) = Co 1 - + 31 5 Since other is an integer maependent soution is . ray orx Jy zo

Post other questions separately

Thank you